The chemist wants 100 gallons when complete. He is mixing a 24% solution with a 50% solution to get a 36% solution. How many gallons of each will he use?
Answer by mananth(16946)

when complete. He is mixing a 24% solution with a 50% solution to get a 36% solution. How many gallons of each will he use?
..
x gallons ------------ 24%
(100-x)gallons ------- 50%
Mixture = 100 gallons--36%
...
0.24x+(100-x)*0.5 = 0.36*100
0.24x+50-0.5x=36
0.24x-0.50x=36-50
-0.26x=-14
/0.26
x= 53.8 gallons of 24%
46.2 gallons of 50%
...
m.ananth@hotmail.ca